Ganymede is one of the moons of Jupiter. Ganymede’s period of revolution around Jupiter is 7.16 Earth days and an orbital radius of about 1.07  m. Find the mass of Jupiter.


1
Expert's answer
2022-02-10T08:36:49-0500

Explanation & Calculation


  • Apply Newton's gravitational law for the planet and its moon.

"\qquad\qquad
\begin{aligned}
\small G\frac{Mm}{r^2}&=\small mr\omega^2\\&=\small mr\frac{4\pi^2}{T^2}\\
\small M&=\small \frac{1}{G}.\frac{4\pi^2.r^3}{T^2}\\
&=\small \frac{1}{6.67\times10^{-11}\,Nm^2kg^{-2}}.\frac{4\pi^2.(1.07\times10^6\times10^3\,m)^3}{(7.16\times24\times3600\,s)^2}\\
&=\small 1.89\times10^{27}\,kg
\end{aligned}"